# Block-Recurrent Dynamics in ViTs

### v1

@@ Appendix A Training Block Recurrent Foundation Models
All Raptor variants ($k=2,3,4$) are trained on top of DINOv2 (ViT-B with registers) (darcet2023vision) and use the same transformer architecture. The model utilizes a feature dimension of $784$, an MLP ratio of $4$, and $12$ multi-head attention heads. The depth-scale MLP consists of a linear layer expanding from dimension 1 to 16, followed by a SiLU activation, and a second linear layer mapping from 16 to $3\times\text{dim}$. This produces three separate scaling vectors. For each Raptor  variant in Table 1 and Figure 7, we train train three models, each with a different seed. We measure the performance of each model and report aggregate statistics. Below, we provide the hyperparameters and training settings used to train Raptor, describing the layer divisions for different values of $k$ followed by the training procedure.
@@ A.3 Linear Probe Fine-tuning
For NYUv2, we adopt an approach similar to oquab2023dinov2. Specifically, we use images at a $480\times 640$ resolution and center pad them so that the dimensions are multiples of 14. We feed the images through the model and extract the predictions from the final layer. The cls token is concatenated with the patch tokens, and the spatial resolution is upsampled by a factor of 4. Both the cls and patch tokens are upscaled, after which the cls token is concatenated to each patch token. We treat this representation as the “logits.” To obtain depth, we normalize the logits with a softmax and compute the weighted average of the centers of 256 evenly spaced bins. Then, we upsample this representation to $480\times 640$ and consider the result our depth. For training, we use the loss function introduced by bhat2021adabins.
@@ Appendix E Levin Complexity under 0-BRH
There is a long line of work linking plain Kolmogorov complexity kolmogorov1965three; chaitin1977algorithmic to its resource-bounded variants, notably Levin’s time-bounded measure, which penalizes programs by both description length and runtime (the $|p|+\log T$ form) (levin1973universal; li2008introduction). Deep learning research has not remained isolated from these ideas: connections have been explored both for discovering simpler neural networks SCHMIDHUBER1997857 and for contextualizing runtime priors schmidhuber2002speed within neural architectures.
In our case, the proof is a direct application of this standard framework as formalized in li2008introduction: once we exhibit a prefix-free program $p$ that computes $\bm{f}_{\ell}$ and bound its running time $T(p)$, Levin complexity immediately yields $K_{\mathrm{Levin}}(\bm{f}_{\ell})\leq|p|+\log T(p)+O(1)$. The only domain-specific ingredients are (i) a BRH-based encoding that prices the tied blocks and schedule at $\sum_{j=1}^{k}DL_{U}(\bm{\theta}(\bm{B}_{j}))+O(k\log\ell)$ using standard self-delimiting integer codes, and (ii) a runtime-parity assumption linking $T(p)$ to the model’s untied runtime $R(\bm{f}_{\ell})$ up to machine and invariance constants. Formally:
@@
We follow the standard time-bounded description framework of levin1973universal, and contextualized it to the 0-BRH setting (perfect reconstruction along the entire depth trajectory), in the spirit of the sketch already outlined in the paper. We fix a universal prefix-free machine $U$ and adopt the same finite-precision arithmetic used by the target model to define $DLU(\cdot)$; this ensures that encoding parameters with length $DLU(\bm{\theta}(\bm{B}_{j}))$ bits suffices to reproduce the exact numerical behavior of the corresponding block under that arithmetic. With this convention, “exactly computes” means bit‑for‑bit equivalence with the reference implementation of $\bm{f}_{\ell}$ (under its inference precision).
@@
Finally, we get the runtime bound to derive the Levin Complexity. On the fixed machine $U$, a single evaluation of block $\bm{B}_{j}$ incurs time at most $c\,R(\bm{B}_{j})$ for a machine dependent constant $c>0$ (standard machine invariance up to constants; cf. the invariance discussion in Kolmogorov/Levin‑style arguments (kolmogorov1965three; levin1973universal)). Executing the schedule therefore costs

### v2

@@ Appendix A Training Block Recurrent Foundation Models
All Raptor variants ($k=2,3,4$) are trained on top of DINOv2 (ViT-B with registers) (Darcet et al., 2023) and use the same transformer architecture. The model utilizes a feature dimension of $784$, an MLP ratio of $4$, and $12$ multi-head attention heads. The depth-scale MLP consists of a linear layer expanding from dimension 1 to 16, followed by a SiLU activation, and a second linear layer mapping from 16 to $3\times\text{dim}$. This produces three separate scaling vectors. For each Raptor  variant in Table 1 and Figure 7, we train train three models, each with a different seed. We measure the performance of each model and report aggregate statistics. Below, we provide the hyperparameters and training settings used to train Raptor, describing the layer divisions for different values of $k$ followed by the training procedure.
@@ A.3 Linear Probe Fine-tuning
For NYUv2, we adopt an approach similar to Oquab et al. (2023). Specifically, we use images at a $480\times 640$ resolution and center pad them so that the dimensions are multiples of 14. We feed the images through the model and extract the predictions from the final layer. The cls token is concatenated with the patch tokens, and the spatial resolution is upsampled by a factor of 4. Both the cls and patch tokens are upscaled, after which the cls token is concatenated to each patch token. We treat this representation as the “logits.” To obtain depth, we normalize the logits with a softmax and compute the weighted average of the centers of 256 evenly spaced bins. Then, we upsample this representation to $480\times 640$ and consider the result our depth. For training, we use the loss function introduced by Bhat et al. (2021).
@@ Appendix E Levin Complexity under 0-BRH
There is a long line of work linking plain Kolmogorov complexity Kolmogorov (1965); Chaitin (1977) to its resource-bounded variants, notably Levin’s time-bounded measure, which penalizes programs by both description length and runtime (the $|p|+\log T$ form) (Levin, 1973; Li et al., 2008). Deep learning research has not remained isolated from these ideas: connections have been explored both for discovering simpler neural networks Schmidhuber (1997) and for contextualizing runtime priors Schmidhuber (2002) within neural architectures.
In our case, the proof is a direct application of this standard framework as formalized in Li et al. (2008): once we exhibit a prefix-free program $p$ that computes $\bm{f}_{\ell}$ and bound its running time $T(p)$, Levin complexity immediately yields $K_{\mathrm{Levin}}(\bm{f}_{\ell})\leq|p|+\log T(p)+O(1)$. The only domain-specific ingredients are (i) a BRH-based encoding that prices the tied blocks and schedule at $\sum_{j=1}^{k}DL_{U}(\bm{\theta}(\bm{B}_{j}))+O(k\log\ell)$ using standard self-delimiting integer codes, and (ii) a runtime-parity assumption linking $T(p)$ to the model’s untied runtime $R(\bm{f}_{\ell})$ up to machine and invariance constants. Formally:
@@
We follow the standard time-bounded description framework of Levin (1973), and contextualized it to the 0-BRH setting (perfect reconstruction along the entire depth trajectory), in the spirit of the sketch already outlined in the paper. We fix a universal prefix-free machine $U$ and adopt the same finite-precision arithmetic used by the target model to define $DLU(\cdot)$; this ensures that encoding parameters with length $DLU(\bm{\theta}(\bm{B}_{j}))$ bits suffices to reproduce the exact numerical behavior of the corresponding block under that arithmetic. With this convention, “exactly computes” means bit‑for‑bit equivalence with the reference implementation of $\bm{f}_{\ell}$ (under its inference precision).
@@
Finally, we get the runtime bound to derive the Levin Complexity. On the fixed machine $U$, a single evaluation of block $\bm{B}_{j}$ incurs time at most $c\,R(\bm{B}_{j})$ for a machine dependent constant $c>0$ (standard machine invariance up to constants; cf. the invariance discussion in Kolmogorov/Levin‑style arguments (Kolmogorov, 1965; Levin, 1973)). Executing the schedule therefore costs